Circular No. 7549

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2000 Y2 B. Skiff, Lowell Observatory, reports the discovery of a comet by the

LONEOS program. Confirming CCD images by L. Wasserman (1.07-m Lowell Observatory telescope) show a coma diameter of ∼ 9′′ and a tail ∼ 14′′ long toward the southwest.

2000 UT α_{2000} δ_{2000} m_{1} Observer
Dec. 27.33505 9^{h}24^{m}06.^{s}52 +0^{o}43′26′′.5 17.2 Skiff
27.35395 9 24 06.78 +0 43 24.3  _{”}
27.37290 9 24 06.98 +0 43 22.2  _{”}
27.39192 9 24 07.18 +0 43 19.7  _{”}
27.42362 9 24 07.41 +0 43 14.0 18.2 Wasserman
27.43262 9 24 07.48 +0 43 12.9  _{”}

SUPERNOVA 2000fp IN ANONYMOUS GALAXY A. Clocchiatti, Pontificia Universidad Catolica de Chile (PUC), on

behalf of the PUC-Padova Supernova Search (cf. IAUC 7537; including P. Arevalo, C. Rada, F. Barrientos, D. Minniti, and F. Courbin, PUC; M. Turatto, G. Altavilla, S. Benetti, E. Cappellaro, and A. Pastorello, Osservatorio di Padova; and B. Leibundgut and J. Spyromilio, European Southern Observatory), reports the discovery of a supernova (V = 22.9) on three images taken on Dec. 17.3 UT with the 2.2-m telescope (+ WFI) at ESO, La Silla. SN 2000fp is located at α = 5^{h}38^{m}01^{s}.27, δ = −23^{o}46′34′′.1 (equinox 2000.0), which is 3′′.2 south and 2′′.9 west of the host galaxy’s nucleus. No object appears at this location on images taken on Nov. 17.24. A spectrum (range 333–747 nm, resolution 1.6 nm) with low signal-to-noise ratio, obtained on Dec. 22.24 with the ESO 3.6-m telescope (+ EFOSC2), shows a relatively blue continuum with few broad absorption features, two of which match Hβ and Hγ once the spectrum is blueshifted to the parent galaxy rest frame (z = 0.30, from several narrow absorption features in a separate spectrum). SN 2000fp appears to be of type II, close to maximum.

COMET 73P/SCHWASSMANN-WACHMANN 3 Additional visual m_{1} estimates for component C (cf. IAUC 7534): Dec.

4.52 UT, 11.6 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 5.84, 10.3 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector); 23.85, 9.9 (Yoshida).

2000 December 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green